OpenStudy (anonymous):

Add and simplify:

OpenStudy (anonymous):

13m^2 --------- 6 13m ------ 12 13m ------- 6

OpenStudy (anonymous):

@E.ali

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Just answer to :5 m +8m=(5+8)m=13mand have /6 too so have : 13m/6

OpenStudy (anonymous):

thx!